OpenStudy (anonymous):

Let v=i+j and B=2i+3j-k, then find the unit, in the direction of a-b?

OpenStudy (zzr0ck3r):

you mean v - b?

OpenStudy (zzr0ck3r):

v = <1,1,0> b = <2,3,-1> v-b = <1-2,1-3,0-(-1)> = <-1,-2,1> unit vector is \(\frac{1}{\sqrt{(-1)^2+(-2)^2+1^2}}<-1,-2,1>= \frac{1}{\sqrt{6}}<-1,-2,1>\\=<\frac{-1}{\sqrt{6}},\frac{-2}{\sqrt{6}},\frac{1}{\sqrt{6}}>\)

OpenStudy (anonymous):

..